# Burlingame, CA 94010, USA

### v3

@@ Abstract
This paper introduces a visual process modeling tool for AI and machine learning in big data contexts, utilizing the LangGraph framework to construct intelligent Spark agents. The tool represents key machine learning stages—data preprocessing, feature engineering, model training, and evaluation—as modular components. Analysts can visually design workflows, which are then automatically translated into optimized Spark code for execution. This approach simplifies the complexity of Apache Spark, reduces the learning curve associated with Scala, and enhances code reusability. The paper discusses the theoretical foundations, key technologies, and evaluates the effectiveness of the proposed solution.

### v4

@@ Abstract
This paper presents a Spark-based modular LangGraph framework, designed to enhance machine learning workflows through scalability, visualization, and intelligent process optimization. At its core, the framework introduces Agent AI, a pivotal innovation that leverages Spark’s distributed computing capabilities and integrates with LangGraph for workflow orchestration.
Agent AI facilitates the automation of data preprocessing, feature engineering, and model evaluation while dynamically interacting with data through Spark SQL and DataFrame agents. Through LangGraph’s graph-structured workflows, the agents execute complex tasks, adapt to new inputs, and provide real-time feedback, ensuring seamless decision-making and execution in distributed environments. This system simplifies machine learning processes by allowing users to visually design workflows, which are then converted into Spark-compatible code for high-performance execution.
The framework also incorporates large language models through the LangChain ecosystem, enhancing interaction with unstructured data and enabling advanced data analysis. Experimental evaluations demonstrate significant improvements in process efficiency and scalability, as well as accurate data-driven decision-making in diverse application scenarios.
This paper emphasizes the integration of Spark with intelligent agents and graph-based workflows to redefine the development and execution of machine learning tasks in big data environments, paving the way for scalable and user-friendly AI solutions.